40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1631     Entry time: Fri May 29 18:57:09 2009
Author: steve 
Type: Update 
Category: PSL 
Subject: the laser is back 

Steve, Rob and Alberto

 

Starting capacitor 216 miroFarad was installed on the compressor. Water lines were connected to the MOPA as corrected, so the flow meter readings are logical.

Now IN means flowing water in the direction of black arrow on the hose.

We struggled with the Neslab presetting:  temp, bauds rate  and other  unknowns  till Rob found the M6000 manual on Peter king's website.

Alberto realized that the chiller temp had to be reset to 20C on water chiller.

I put 1mg of Chloramin T into the water to restrict the growth of algae in the bath.

The NPRO heat sink  was around ~20C without flow meter wheel rotation  and the PA body ~25C by touch of a finger

I just opened up the needle valve a litle bit so the flow meter wheel would started rotating slowly.

That small glitch at the end of this 3 hrs plot shows this adjustment.

Attachment 1: laserisback.jpg  90 kB  | Hide | Hide all
laserisback.jpg
ELOG V3.1.3-